OpenStudy (anonymous):

How do you factor this polynomial? Ex. f(x)=x^3-x^2+16x-16

OpenStudy (anonymous):

Factor by grouping: make f(x)=0 (x^3-x^2)+(16x-16) = 0 Find the GCF of each group and take it outside the parenthesis: x^2(x-1)+16(x-1)=0 Factors: (x^2+16)(x-1)=0 Use the Zero Factor Property to find x terms: ab=0, if a=0 or b=0 x^2+16=0 x-1=0 x^2=-16 x = 1 Do the square root of both sides \[x=\pm \sqrt{-16}\] \[x=\pm4i\]

OpenStudy (anonymous):

Do you have any questions?

OpenStudy (anonymous):

No I understand it, thank you.